# Phonon-driven multipolar dynamics in a spin-orbit coupled Mott insulator

### v3

@@ Abstract
Figure 2: Dynamics of $\langle\tau_{y}\rangle$ in the presence of phase coherent two-phonon drive with weak dissipation (see text for details). (a) In the paramagnet $T\!=\!1.1T_{\rm oct}$, with drive strength $A\!=\!0.05$ (resonant phonon amplitude $Q_{0}\!\approx\!0.5$) and various phase differences $\phi\in[-\pi/2,\pi/2]$, showing that two-phonon drive can be used ‘train’ the sign of octupolar order. (b) In the ferro-octupolar ordered phase $T\!=\!0.8T_{\rm oct}$, with a larger drive strength $A\!=\!0.2$ (resonant phonon amplitude $Q_{0}\!\approx\!1.5$) showing that with a suitable choice of phase $\phi\!=\!\pi/2$ it can be used to switch from positive to negative octupolar order.
@@
Dynamical switching of octupolar order.— We next ask if, using a strong enough phonon drive, this same mechanism can also be used to switch the sign of octupolar order for $T\!<\!T_{\rm oct}$. Fig. 2(b) shows the impact of the two-phonon coherent drive for $\phi=\pi/2$ at $T/T_{\rm oct}=0.8$. Using a drive which leads to a resonant phonon amplitude $Q_{0}\sim 1.5$, we find that octupolar order in the positive domain switches sign, while the octupolar order in the negative domain is nearly unaffected. The slow evolution of $\langle\tau_{y}\rangle$ at late times, and the nearly unchanged $\langle\tau_{y}\rangle$ in the ‘--’ domain, indicate that heating effects are weak at these timescales. In order for this switching to occur, the phonon-induced field $(-\lambda^{2}Q_{0}^{2}/\Omega)\sin\phi$ must overcome the octupolar Weiss field in the ordered phase. We find that the phonon-induced field is too small to flip pseudospins deep inside a domain, but it is able to induce switching at domain walls where the Weiss field is strongly suppressed. Thus, negative domains must grow by flipping pseudospins at the boundary until the entire octupolar $\langle\tau_{y}\rangle$ flips. In Fig.3, we depict the time evolution of pseudospins under a two-phonon drive starting with an initial 3D domain wall configuration and a thermal population of phonon momenta. We focus on a single 2D slice through the lattice, with colours indicating positive $\langle\tau_{y}\rangle$ (red) and negative $\langle\tau_{y}\rangle$ (blue) and the black sites denoting the non-magnetic sublattice. One can clearly see the two-phonon drive inducing a shrinking of the $+\langle\tau_{y}\rangle$ domain as a function of time via spin flipping at the domain walls. In the bottom right panel of Fig.3 we show the expectation values of $\tau_{x},\tau_{y}$ and $\tau_{z}$ averaged over the lattice; we attribute the rapid oscillations to the high-frequency phonons and the apparent $x$-translational symmetry breaking to the initial distribution of phonon momenta. We provide similar pictures of the lattice spins for the switching presented in Fig. 2(b) in the SM [72], as well as a simple theoretical model of this mechanism.
@@
Figure 3: Visualization of octupolar order switching starting with an initial domain wall configuration of pseudospins. We drive the $Q_{x}$ and $Q_{z}$ phonons with a relative phase of $\pi/2$ and show three time points around $t\sim\!7\hbar/J_{0}$, the point at which switching begins to occur. Here, we depict spins having $\langle\tau_{y}\rangle>0$ as red squares and $\langle\tau_{y}\rangle<0$ as blue squares. We initialize half the spins with $\langle\tau_{y}\rangle=+1.0$ and half with $\langle\tau_{y}\rangle=-1.0$. As time progresses, one can clearly see the shrinking of the positive (red) domain and growth of the negative (blue) domain as reflected in $\langle\tau_{y}\rangle$ averaged over the whole lattice, which is shown in the bottom right panel. Note that the rapid oscillation in the $\tau_{x},\tau_{y}$ and $\tau_{z}$ magnetization appears at the phonon frequency, while the apparent translational symmetry breaking is due to the initial phonon momenta, which are chosen from a thermal distribution.
@@
While experimental verification of our theoretical results on training and switching of octupolar order switching would necessitate a method of detecting the presence of octupolar order, this has proven notoriously difficult to do using conventional experimental techniques. Remarkably, Ref.[75] presents a promising route by which to unambiguously diagnose both the presence and, significantly, sign of the octupolar order parameter. Explicitly, rotational anisotropy second harmonic generation (SHG) measurements are conducted and the presence of octupolar order is characterized through interference between the time reversal even and odd channels, with the sign of the octupolar order determining the positions of the measured intensity maxima. Thus, we propose a tr-SHG experiment with a first measurement conducted at $t=0$ and a second conducted at time $t\sim 25\hbar/J_{0}$ ($\sim 16\rm ps$) looking for changes in the position of intensity maxima as a potential route to diagnose the training and switching of the octupolar order parameter. In the SM [72], we also provide further plots showing how our results depend on the phenomenological damping parameter $\eta$.
Summary.— We have employed Monte Carlo and molecular dynamics simulations capable of modelling lattice degrees of freedom, in the form of local Einstein phonons, in addition to multipolar degrees of freedom to study Mott insulators which host non-Kramers doublets. Our equilibrium phase diagram for the model hosts ferro-octupolar or antiferro-quadrupolar ordered phases at low temperature, as partly described by previous work, but shows that the quadrupolar order is accompanied by lattice displacements. We next studied the out-of-equilibrium multipolar-phonon dynamics in the presence of a uniform excitation of Einstein phonon modes. Emulating pump-probe type experiments by giving a uniform impulsive kick to one phonon mode, we observed a characteristic response in the non-excited phonon below the octupolar ordering temperature which is absent in the disordered phase, leading to a dynamical signature of equilibrium octupolar order. Secondly, we model a continuous excitation of both phonons, with a relative phase between them, and observe the generation of octupolar order consistent with an effective field predicted by a lab frame Magnus expansion. This also led us to a proposed route to ultrafast picosecond timescale switching of octupolar order using phonons. Our work opens up unique directions in the study of Mott insulators playing host to multipolar moments, including potential routes by which to detect and control complex hidden orders.
@@
This research was funded by an NSERC Discovery Grant (AP), an Ontario Graduate Scholarship (KH), and an NSERC CGS-D fellowship (RS). We thank Sreekar Voleti for initial collaboration on this project. We acknowledge Dave Hsieh, Emily Zhang, William Bateman-Hemphill, and Cristian Batista for extremely helpful discussions. Numerical computations were performed on the Niagara supercomputer at the SciNet HPC Consortium and the Digital Research Alliance of Canada.

### v4

@@ Abstract
Figure 2: Dynamics of $\langle\tau_{y}\rangle$ in the presence of short-time phase coherent two-mode phonon drive with weak dissipation. The drive $H_{\rm drive}^{\rm ph}$ is turned on from $t=0\hbar/J_{0}$ to $t=3\hbar/J_{0}$ and then turned off. (a) In the paramagnet $T\!=\!1.1T_{\rm oct}$, with drive strength $A\!=\!0.5$ (peak phonon amplitude $Q_{0}\!\approx\!0.75$) and various phase differences $\phi\in[-\pi/2,\pi/2]$, showing that driven phonons can select the sign of generated octupolar order. (b) In the ferro-octupolar ordered phase $T\!=\!0.8T_{\rm oct}$, with varying drive strength $A$, which induces a peak phonon amplitude $Q_{0}$ between $2-3$. With a suitable phase choice $\phi\!=\!\pi/2$, it can be used to decrease the octupolar order parameter $\langle\tau_{y}\rangle$, and even switch from positive to negative. Inset: Phonon coordinate $Q_{x}(t)$ for $A\!=\!2$, showing its amplitude build-up and eventual dissipation after the drive is switched off.
Dynamical switching of octupolar order.— We next ask if this pulsed phonon drive can also be used to switch the sign of octupolar order for $T\!<\!T_{\rm oct}$. Fig. 2(b) shows the impact of the two-phonon coherent drive for $\phi=\pi/2$ at $T/T_{\rm oct}=0.8$. Using a drive which leads to a peak phonon amplitude $Q_{0}\!\sim\!1$-$3$ (i.e., $\sim\!1$-$3\%$ atomic displacement), we find that octupolar order in the positive domain is suppressed or can even switch sign, depending on the drive amplitude while the octupolar order in negative domains (not shown) is nearly unaffected. Furthermore, the slow evolution of $\langle\tau_{y}\rangle$ at late times indicate that heating effect of pseudospins is weak at these timescales, although additional cooling might be needed in experiments to extract the energy dissipated by the phonon modes. Nonlinear phononics experiments on CoF2 have realized large atomic displacements $\sim 2\%$ of Raman-active modes by driving an infrared-active phonon with $500$ fs pulse of 10MV/cm electric field [15]; such methods could potentially be extended to the osmates explored in this work.
@@
We find that octupolar switching is mediated by flipping of pseudospins at domain wall boundaries by the phonon-induced field $(-\lambda^{2}Q_{0}^{2}/\Omega)\sin\phi$; for $\phi\!=\!\pi/2$, this promotes growth of $\langle\tau_{y}\rangle\!<\!0$ domains. In Fig.3, we depict this time evolution of pseudospins under a two-phonon drive starting with an initial 3D domain wall configuration with periodic boundary conditions, and a thermal population of phonon momenta. We show a single 2D slice through the lattice, with colours indicating positive $\langle\tau_{y}\rangle$ (red) and negative $\langle\tau_{y}\rangle$ (blue) and the black sites denoting the non-magnetic sublattice. One can clearly see the two-phonon drive inducing a shrinking of the $+\langle\tau_{y}\rangle$ domain as a function of time. The bottom right panel of Fig.3 shows $\langle\tau_{y}\rangle$ averaged over the full lattice, evincing growth of octupolar magnetization; we attribute the superposed rapid oscillations to the high-frequency phonons. Additional data and an effective model of this mechanism is given in the SM [80].
@@
Figure 3: Visualization of octupolar order switching for an initial two-domain configuration (assuming periodic boundary conditions). We drive the $Q_{x}$ and $Q_{z}$ phonons with a relative phase $\phi\!=\!\pi/2$ for a short time $t=3\hbar/J_{0}$ ($\sim 1$ ps) and show configurations at representative times. Red and blue squares respectively depict pseudospins with $\langle\tau_{y}\rangle>0$ and $\langle\tau_{y}\rangle<0$. We clearly see the shrinking of the positive (red) domain and growth of the negative (blue) domain. This domain evolution is also reflected in the net octupolar magnetization $\langle\tau_{y}\rangle$ averaged over the whole lattice (bottom right panel) which is initially zero but evolves to negative values.
@@
Recent remarkable experiments [85] have shown that one can unambiguously diagnose the presence and, significantly, sign of ferro-octupolar order using rotational anisotropy second harmonic generation (SHG). The octupolar order is characterized through interference between the time reversal even and odd channels, with the sign of the octupolar order determining the positions of the measured intensity maxima [85]. We propose a tr-SHG experiment which compares the position of intensity maxima of equilibrium versus the driven system at time $t\sim 3.0\hbar/J_{0}$ ($\sim 2.0\rm ps$) to diagnose the theoretically predicted build-up and switching of octupolar order. Additional plots showing dependence of our results on the phonon damping parameter $\eta$ are given in the SM [80].
Summary.— We have developed computational tools to simulate the equilibrium phases and nonequilibrium dynamics of multipolar degrees of freedom coupled to dissipative Einstein phonons in multipolar Mott insulators. Our equilibrium phase diagram for the model hosts ferro-octupolar and quadrupolar orders at low temperature, as partly described by previous work, but additionally shows that the quadrupolar order is accompanied by weak lattice displacements. We have also shown how pump-probe techniques and two-mode phonon drive can uncover unique dynamical signatures and enable the ultrafast control of the ferro-octupolar state. This study opens up distinct directions in the detection and control of complex hidden orders in quantum materials.
@@
This research was funded by an NSERC Discovery Grant (AP), an Ontario Graduate Scholarship (KH), and an NSERC CGS-D fellowship (RS). We thank Sreekar Voleti, David Hsieh, Prashant Padmanabhan, Emily Zhang, William Bateman-Hemphill, and Cristian Batista for extremely helpful discussions. Numerical computations were performed on the Niagara supercomputer at the SciNet HPC Consortium and the Digital Research Alliance of Canada. Data for the figures in this work is available, see [86].